a body is projected from the bottom of smooth inclined plane with a velocity of 20m/sif it is just sufficient to carry it to a top in 4s find the inclination and height of the plane

Grade:12

1 Answers

Chandan
6 years ago
let inclination be x
if time = 4 sec
20=mgsinx*4   (v=u+at)
mgsinx=5
l=80-8mgsinx    (s=ut+1\at2)
l=40
conserving energy
200=10lsinx
x=300
hope it helps please approve if it does